OpenStudy (emilylovesgiraffes):

Given the following chemical equation, how many moles of carbon are necessary to react with 5.52 moles of SO2?5C(s) + 2SO2 (g) = CS2 (l) + 4CO (g)

OpenStudy (danjs):

start with the given 5.52 mol SO2. Use mole ratio from the balanced equation to convert to Mol C

OpenStudy (danjs):

\[\large (\frac{ 5.52~mol~SO2 }{ 1 })*(\frac{ 5~mol~C }{ 2~mol~SO2 })=\]